# Sylvester domains and pro-$p$ groups

### v1

@@ Abstract.
Let $G$ be a finitely generated torsion-free pro-$p$ group containing an open free-by-$\mathbb{Z}_{p}$ pro-$p$ subgroup. We show that the completed group algebra $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain. Moreover the inner rank $\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}(A)$ of a matrix $A$ over $\mathbb{F}_{p}[\![G]\!]$ can be calculated by approximation by ranks corresponding to finite quotients of $G$, that is, if $G=G_{1}>G_{2}>\ldots$ is a chain of normal open subgroups of $G$ with trivial intersection and $A_{i}$ is the matrix over $\mathbb{F}_{p}[G/G_{i}]$ obtained from the matrix $A$ by applying the natural homomorphism $\mathbb{F}_{p}[\![G]\!]\to\mathbb{F}_{p}[G/G_{i}]$, then $\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}(A)=\displaystyle\lim_{i\to\infty}\frac{\operatorname{rk}_{\mathbb{F}_{p}}(A_{i})}{|G:G_{i}|}$. As a consequence, we obtain a particular case of the mod $p$ Lück approximation for abstract finitely generated subgroups of free-by-$\mathbb{Z}_{p}$ pro-$p$ groups.
@@ 1. Introduction
A Sylvester rank function on a ring $R$ is a function taking non-negative real values on matrices over $R$ and satisfying a series of conditions (see Paragraph 2.4) resembling the conditions of the rank function of matrices over a field. The value of a Sylvester rank function $\operatorname{rk}$ on a matrix $A$ over $R$ is bounded from above by its inner rank $\operatorname{rk}(A)\leq\operatorname{irk}_{R}(A)$. If $\operatorname{irk}_{R}$ is itself a Sylvester rank function, then the ring $R$ is called a Sylvester domain. This notion appeared first implicitly in the works of P. Cohn and was explicitly defined by W. Dicks and E. Sontag [DS78]. One remarkable property of a Sylvester domain $R$ is that it has a universal division ring of fractions $\mathcal{D}_{R}$ such that for every matrix $A$ over $R$, its rank over $\mathcal{D}_{R}$ is equal to its inner rank $\operatorname{irk}_{R}(A)$.
@@
In the case $R=K[\Gamma]$ is a group algebra of a free group $\Gamma$ over a field $K$, it was proven by P. Cohn [Coh06, Theorem 7.11.8, Proposition 5.5.1] that $K[\Gamma]$ is a Sylvester domain. In [HLA22], F. Henneke and D. López-Álvarez considered the case where $\Gamma$ is a $-\mathbb{Z}$ group and showed that $K[\Gamma]$ is a Sylvester domain if and only if every left finitely generated projective $K[\Gamma]$-module is free.
@@
We would like to notice that, for an abstract group $\Gamma$, if $K$ is a field of characteristic 0 and $K[\Gamma]$ is a Sylvester domain, then $\operatorname{irk}_{K[\Gamma]}$ coincides with the von Neumann Sylvester rank function $\operatorname{rk}_{\Gamma}$ (for the definition of $\operatorname{rk}_{\Gamma}$, see [JZ19b, Section 2]).
@@
In this paper we want to understand for which pro-$p$ groups $G$ the completed group algebra $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain. Often, the results in combinatorial group theory and combinatorial pro-$p$ group theory run in parallel; although their proofs are quite different. For example, the theorem of J. Stallings ([Sta68]) saying that a torsion-free and virtually free group is itself free was inspired by the previous and analogous result on free pro-$p$ groups due to J.-P. Serre in [Ser65]. Later in [Lub82], A. Lubotzky described how the classical theorems of M. Hall, L. Greenberg and A. Howson about finitely generated subgroups of free groups all have analogues for free pro-$p$ groups. More recently, a new proof of the Hanna Neumann Conjecture (first proven independently by J. Friedman and I. Mineyev in 2011) was given by the first author in [Jai17] using homological methods that apply both to the abstract and pro-$p$ versions of the conjecture. This result was later extended to non-solvable pro-$p$ Demushkin groups and their discrete counterparts, the hyperbolic surface groups, in [JZS19] and [AJ22] respectively. In the case of a pro-$p$ group $G$ there exists a Sylvester function on $\mathbb{F}_{p}[\![G]\!]$, which is an analogue of the von Neumann Sylvester rank function and we also denote it by $\operatorname{rk}_{G}$ (see Example 2.5 for the definition). Its definition resembles the Lück approximation [Lüc94, JZ19b].
@@
If $F$ is a finitely generated free pro-$p$ group, then $\mathbb{F}_{p}[\![F]\!]$ is isomorphic to a ring of non-commutative formal power series over $\mathbb{F}_{p}$, and so from [Coh74] we know that $\mathbb{F}_{p}[\![F]\!]$ is a Sylvester domain. In [JZ20], the first author proved that in this case $\operatorname{irk}_{\mathbb{F}_{p}[\![F]\!]}=\operatorname{rk}_{F}$. In this paper we extend this result to finitely generated free-by-$\mathbb{Z}_{p}$ pro-$p$ groups.
@@
We do not know whether the last conclusion of the theorem holds always if $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain. We leave this as a question.
@@
One striking consequence of Theorem 1.1 is that $\operatorname{rk}_{G}$ takes only integer values. This provides new cases for which the pro-$p$ Atiyah conjecture holds (see [JZ19b, Conj. 11.1]).
@@
Recall that the usual Atiyah conjecture for torsion-free abstract groups $\Gamma$ states that $\operatorname{rk}_{\Gamma}$ takes only integer values (see [JZ19b] for more details). Notice that the pro-$p$ Atiyah conjecture implies also pro-$p$ Kaplansky conjecture: $\mathbb{F}_{p}[\![G]\!]$ does not have non-trivial zero divisors if $G$ is torsion-free. Previously the pro-$p$ Atiyah conjecture was only known for pro-$p$ groups which are residually-(torsion-free $p$-adic analytic) ([JZS19, Cor. 5.5]). This class includes all infinite Demushkin groups, which are the pro-$p$ groups that satisfy Poincaré duality in dimension $2$. Infinite Demushkin $G$ groups comprise an important class of free-by-$\mathbb{Z}_{p}$ pro-$p$ groups, and while the Atiyah conjecture for them was previously known, the statement that $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain is new. In Paragraph 2.2 we present examples of finitely generated torsion-free pro-$p$ groups containing an open free-by-$\mathbb{Z}_{p}$ pro-$p$ subgroup but which are not free-by-$\mathbb{Z}_{p}$ themselves.
Let us describe the structure of the paper and the main ideas behind the proof of Theorem 1.1 comparing it with the proof of the main result in [HLA22]. We let $d(G)$ denote the cardinality of a minimal set of topological generators of a pro-$p$ group $G$.
On one hand, all projective $\mathbb{F}_{p}[\![G]\!]$-modules are free for the completed group algebra $\mathbb{F}_{p}[\![G]\!]$, and so, in contrast to the case of abstract free-by-cyclic groups considered in [HLA22], there is no difference for $\mathbb{F}_{p}[\![G]\!]$ to be a pseudo-Sylvester domain or a Sylvester domain.
@@
For an arbitrary finitely generated virtually free-by-$\mathbb{Z}_{p}$ pro-$p$ group $G$, this construction provides an embedding of $\mathbb{F}_{p}[\![G]\!]$ in an Artinian ring $\mathcal{Q}$. The next step is to show that under this embedding all full matrices over $\mathbb{F}_{p}[\![G]\!]$ become invertible over $\mathcal{Q}$. The main tool to prove this is Theorem 2.8 proven in [JZ20]. We achieve this in Section 6. This shows that $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain.
@@
We finish the paper with an appendix where we show that if $R$ is a profinite ring and $R_{\Sigma}$ is a localization of $R$ with respect to a collection $\Sigma$ of a square matrices over $R$, then $R_{\Sigma}$ admits a Hausdorff ring topology such that the map $R\to R_{\Sigma}$ is continuous. This is used only in the proof of Claim 3.2 (see also the remark afterwards).
@@ Acknowledgments
This work is partially supported by the grant PID2020-114032GB-I00 of the Ministry of Science and Innovation of Spain and by the ICMAT Severo Ochoa project CEX2019-000904-S4. We would like to thank Alan Logan and Henry Wilton for pointing out several examples of torsion-free virtually $-\mathbb{Z}$ but not $-\mathbb{Z}$ groups.
@@ 2.1. General notation
If $G$ is a pro-$p$ group, we denote the Frattini subgroup of $G$ by $\Phi(G)=G^{p}[G,G]$ and the lower $\bm{p}$-central series of $G$ by $G_{1}=G$ and $G_{i}=G_{i-1}^{p}[G_{i-1},G]$, so that $G_{2}=\Phi(G)$. We stack the commutators $[a,b]=a^{-1}b^{-1}ab$ on the right:
$[a_{1},a_{2},a_{3},\ldots,a_{n}]=[[\cdots[[a_{1},a_{2}],a_{3}],\cdots],a_{n}]\,.$
@@
When we say that a pro-$p$ group $G$ is finitely generated, or that $X$ is a generating set for $G$, it will always mean generation in the topological sense. We recall that if $w_{0}(G)$ denotes the smallest cardinality of a fundamental system of neighborhoods of $1$ in $G$ – it’s local weight – then $w_{0}(G)=\max\{d(G),|\mathbb{N}|\}$ ([RZ10, Cor. 2.6.3]). In particular, every closed subgroup of a finitely generated pro-$p$ group is at most countably generated.
@@ 2.2. Torsion-free virtually free-by-$\mathbb{Z}_{p}$ groups
There are many abstract torsion-free groups which are virtually $-\mathbb{Z}$ but not $-\mathbb{Z}$ themselves. One explicit family of such groups is given by the presentations $\Gamma_{e}=\langle a,b,x,y\mid[a,b]^{e}=[x,y]^{e}\rangle$ for $e>1$ (see [BFMT09, Example 7.2]). More examples arise amongst 3-manifold groups such as fundamental groups of knot complements with Alexander polynomial $1$. One family of such knots is given by the Pretzel links $L(2m+1,2n+1,2q+1)$ satisfying $(m+n+1)(m+p+1)=m(m+1)$, such as $L(-3,5,7)$. The condition on the Alexander polynomial implies that $[G,G]$ is perfect, and such groups virtually fibre over $\mathbb{Z}$ by being virtually special as a combined consequence of the works of I. Agol, P. Przytycki and D. Wise([Ago08, Ago13, PW18]).
In the pro-$p$ case we can construct examples of torsion-free pro-$p$ groups which are virtually free-by-$\mathbb{Z}_{p}$ but not free-by-$\mathbb{Z}_{p}$ themselves using the ideas of [BFMT09].
@@
Proof.
Write $G$ as the proper free pro-$p$ product with amalgamation $G=A\amalg_{C}B$ of the free pro-$p$ groups $A=F(a_{1},a_{2})$, $B=F(b_{1},b_{2})$ and cyclic subgroup $C=\langle[a_{1},a_{2}]^{p}\rangle\simeq\langle[b_{1},b_{2}]^{p}\rangle$ ([RZ10, Exer. 9.2.6a)]). Since both $A$ and $B$ are torsion-free, $G$ is also torsion-free by [RZ00, Thm. 4.2(b)].
Let $F=F(x_{1},x_{2})$ be another free pro-$p$ group on two generators and consider the surjective homomorphism $\varphi\colon G\to F$ defined by $\varphi(a_{i})=\varphi(b_{i})=x_{i}$ for $i=1,2$.
Claim 2.2.
@@
Let $V=\Phi(F)$. Then, it is clear that $[x_{1},x_{2}]$ is not contained in $[V,V]$ and hence is a non-trivial element of $M=V/[V,V]\simeq\mathbb{Z}_{p}^{n}$. Observe that $[x_{1},x_{2}]+[V,V]$ has finitely many $F$-conjugates which are linearly independent modulo $[V,V]$, and so one can find a homomorphism $\psi_{0}\colon M\to\mathbb{Z}_{p}$ such that $\psi_{0}(x^{-1}[x_{1},x_{2}]x+[V,V])=1$ for all $x\in F$. The desired $\psi$ is then the composition of the natural projection $V\to M$ with $\psi_{0}$. ∎
@@
Let $(\mathcal{G},X)$ be the graph of pro-$p$ groups consisting of two vertices $A$ and $B$ and an edge $C$ connecting them such that $\pi_{1}(\mathcal{G},X)\simeq G$, and let $Y$ be the standard $p$-tree associated with $(\mathcal{G},X)$ ([RZ00, Sec. 4]). We recall that:
@@
Let $Z=K\backslash Y$ and $\mathcal{Z}$ be the associated graph of pro-$p$ groups on $Z$. The edge stabilizers of the $K$ action on $Y$ are the intersections $K\cap gCg^{-1}$ for $g\in G$, which are trivial by construction. Since the vertex stabilizers are all free pro-$p$, the fundamental group $\pi_{1}(\mathcal{Z},Z)$ is also free pro-$p$.
@@
Given that $[a_{1},a_{2}]=[b_{1},b_{2}]$ is an element of both $U\cap A$ and $U\cap B$ that is not trivial under $\theta$, $V(Z)$ is finite (and in fact consists of two elements).
Note that $C$ is contained in $U$, and since it generates the image of $\theta$ we have $U\simeq K\rtimes C$. Also observe that by the construction of $\theta$ and $K$ we have $[G,U]\leq K$. Combining this information with the homeomorphism $K\backslash G\simeq K\backslash U\times U\backslash G$, we get that the induced right action of $C$ in $K\backslash U\times U\backslash G$ is given by
@@
and $Z$ is finite. From [Rib17, Thm. 6.6.1], we conclude that the induced map $\pi_{1}(\mathcal{Z},Z)\to K$ is an isomorphism. ∎
Lastly, observe that $G$ itself cannot be free-by-$\mathbb{Z}_{p}$. By [KM93, Corollary 3.5], $\Gamma$ is residually-$p$, and so it embeds into $G$. Since $[a_{1},a_{2}]$ and $[b_{1},b_{2}]$ are distinct elements in $[\Gamma,\Gamma]\subseteq[G,G]$ with the same $p$-th power, the group $[G,G]$ cannot be free pro-$p$. ∎

### v2

@@ Abstract.
Let $G$ be a finitely generated torsion-free pro-$p$ group containing an open free-by-$\mathbb{Z}_{p}$ pro-$p$ subgroup. We show that the completed group algebra $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain. Moreover the inner rank $\operatorname{irk}_{\mathbb{F}_{p}[\![G]\!]}(A)$ of a matrix $A$ over $\mathbb{F}_{p}[\![G]\!]$ can be calculated by approximation by ranks corresponding to finite quotients of $G$. As a consequence, we obtain a particular case of the mod $p$ Lück approximation for abstract finitely generated subgroups of free-by-$\mathbb{Z}_{p}$ pro-$p$ groups.
@@ 1. Introduction
A Sylvester rank function on a ring $R$ is a function taking non-negative real values on matrices over $R$ and satisfying a series of conditions (see Paragraph 2.4) resembling the conditions of the rank function of matrices over a field. The value of a Sylvester rank function $\operatorname{rk}$ on a matrix $A$ over $R$ is bounded from above by its inner rank $\operatorname{rk}(A)\leq\operatorname{irk}_{R}(A)$. If $\operatorname{irk}_{R}$ is itself a Sylvester rank function, then the ring $R$ is called a Sylvester domain. This notion appeared first implicitly in the works of P. Cohn and was explicitly defined by W. Dicks and E. Sontag [11]. One remarkable property of a Sylvester domain $R$ is that it has a universal division ring of fractions $\mathcal{D}_{R}$ into which it embeds and such that for every matrix $A$ over $R$, its rank over $\mathcal{D}_{R}$ is equal to its inner rank $\operatorname{irk}_{R}(A)$. In particular, $R$ has no zero divisors, see [11, Sec. 0].
@@
In the case $R=K[\Gamma]$ is a group algebra of a free group $\Gamma$ over a field $K$, it was proven by P. Cohn [10, Theorem 7.11.8, Proposition 5.5.1] that $K[\Gamma]$ is a Sylvester domain. In [17], F. Henneke and D. López-Álvarez considered the case where $\Gamma$ is a free-by-$\mathbb{Z}$ group and showed that $K[\Gamma]$ is a Sylvester domain if and only if every left finitely generated projective $K[\Gamma]$-module is free.
@@
We would like to notice that, for an abstract group $\Gamma$, if $K$ is a field of characteristic 0 and $K[\Gamma]$ is a Sylvester domain, then $\operatorname{irk}_{K[\Gamma]}$ coincides with the von Neumann Sylvester rank function $\operatorname{rk}_{\Gamma}$ (for the definition of $\operatorname{rk}_{\Gamma}$, see [22, Section 2]).
@@
In this paper we want to understand for which pro-$p$ groups $G$ the completed group algebra $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain. Often, the results in combinatorial group theory and combinatorial pro-$p$ group theory run in parallel; although their proofs are quite different. For example, the theorem of J. Stallings ([49]) saying that a torsion-free and virtually free group is itself free was inspired by the previous and analogous result on free pro-$p$ groups due to J.-P. Serre in [47]. Later in [32], A. Lubotzky described how the classical theorems of M. Hall, L. Greenberg and A. Howson about finitely generated subgroups of free groups all have analogues for free pro-$p$ groups. More recently, a new proof of the Hanna Neumann Conjecture (first proven independently by J. Friedman and I. Mineyev in 2011) was given by the first author in [21] using homological methods that apply both to the abstract and pro-$p$ versions of the conjecture. This result was later extended to non-solvable pro-$p$ Demushkin groups and their discrete counterparts, the hyperbolic surface groups, in [20] and [3] respectively. Still in the theme of Demushkin groups, in [48] M. Shusterman and P. Zalesskii extended the pro-$p$ version of Howson’s theorem to non-solvable Demushkin groups, for which they also proved the virtual retractions property that was stablished for surface groups by P. Scott in [46].
In the case of a pro-$p$ group $G$ there exists a Sylvester function on $\mathbb{F}_{p}[\![G]\!]$, which is an analogue of the von Neumann Sylvester rank function and we also denote it by $\operatorname{rk}_{G}$ (see Example 2.5 for the definition). Its definition resembles the Lück approximation [33, 22].
@@
If $F$ is a finitely generated free pro-$p$ group, then $\mathbb{F}_{p}[\![F]\!]$ is isomorphic to a ring of non-commutative formal power series over $\mathbb{F}_{p}$, and so from [9] we know that $\mathbb{F}_{p}[\![F]\!]$ is a Sylvester domain. In [24], the first author proved that in this case $\operatorname{irk}_{\mathbb{F}_{p}[\![F]\!]}=\operatorname{rk}_{F}$. In this paper we extend this result to finitely generated free-by-$\mathbb{Z}_{p}$ pro-$p$ groups.
@@
Recall that the pro-$p$ Atiyah conjecture predicts that if $G$ is a pro-$p$ group with the exponent of its torsion elements bounded by $p^{n}$ then $\operatorname{rk}_{G}$ takes values in $p^{-n}\mathbb{Z}$ (see [22, Conj. 11.1]). The following consequence of Theorem 1.1 provides new cases where this conjecture holds.
@@
Notice that the pro-$p$ Atiyah conjecture implies also pro-$p$ Kaplansky conjecture: $\mathbb{F}_{p}[\![G]\!]$ does not have non-trivial zero divisors if $G$ is torsion-free. Previously the pro-$p$ Atiyah conjecture was only known for pro-$p$ groups which are residually-(torsion-free $p$-adic analytic) ([20, Cor. 5.5]). This class includes all infinite Demushkin groups, which are the pro-$p$ groups that satisfy Poincaré duality in dimension $2$. Infinite Demushkin $G$ groups comprise an important class of free-by-$\mathbb{Z}_{p}$ pro-$p$ groups, and while the Atiyah conjecture for them was previously known, the statement that $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain is new. In Paragraph 2.2 we present examples of finitely generated torsion-free pro-$p$ groups containing an open free-by-$\mathbb{Z}_{p}$ pro-$p$ subgroup but which are not free-by-$\mathbb{Z}_{p}$ themselves. Observe that the free kernel is abelian if and only if $G$ is a $p$-adic analytic group (see Remark 2.10 for a full characterization of the $p$-adic analytic groups $G$ for which $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain).
Let us describe the structure of the paper and the main ideas behind the proof of Theorem 1.1 comparing it with the proof of the main result in [17]. We let $d(G)$ denote the cardinality of a minimal set of topological generators of a pro-$p$ group $G$.
On one hand, all projective $\mathbb{F}_{p}[\![G]\!]$-modules are free for the completed group algebra $\mathbb{F}_{p}[\![G]\!]$, and so, in contrast to the case of abstract free-by-cyclic groups considered in [17], there is no difference for $\mathbb{F}_{p}[\![G]\!]$ to be a pseudo-Sylvester domain or a Sylvester domain.
@@
For an arbitrary finitely generated virtually free-by-$\mathbb{Z}_{p}$ pro-$p$ group $G$, this construction provides an embedding of $\mathbb{F}_{p}[\![G]\!]$ in an Artinian ring $\mathcal{Q}$. The next step is to show that under this embedding all full matrices over $\mathbb{F}_{p}[\![G]\!]$ become invertible over $\mathcal{Q}$. The main tool to prove this is Theorem 2.8 proven in [24]. We achieve this in Section 6. This shows that $\mathbb{F}_{p}[\![G]\!]$ is a Sylvester domain.
@@
We finish the paper with an appendix where we show that if $R$ is a profinite ring and $R_{\Sigma}$ is a localization of $R$ with respect to a collection $\Sigma$ of a square matrices over $R$, then $R_{\Sigma}$ admits a Hausdorff ring topology such that the map $R\to R_{\Sigma}$ is continuous. This is used only in the proof of Lemma 3.2 (see also the remark afterwards).
@@ Acknowledgments
This work is partially supported by the grant PID2020-114032GB-I00 of the Ministry of Science and Innovation of Spain and by the ICMAT Severo Ochoa project CEX2019-000904-S4. We would like to thank Alan Logan and Henry Wilton for pointing out several examples of torsion-free virtually free-by-$\mathbb{Z}$ but not free-by-$\mathbb{Z}$ groups. The authors also extend their gratitude to the anonymous reviewer for their valuable comments.
@@ 2.1. General notation
If $G$ is a pro-$p$ group, we denote the Frattini subgroup of $G$ by $\Phi(G)=G^{p}[G,G]$ and the lower $\bm{p}$-central series of $G$ by $G_{1}=G$ and $G_{i}=G_{i-1}^{p}[G_{i-1},G]$, so that $G_{2}=\Phi(G)$. We stack the commutators $a,b]=a^{-1}b^{-1}ab$ on the right:
$a_{1},a_{2},a_{3},\ldots,a_{n}]=[[\cdots[[a_{1},a_{2}],a_{3}],\cdots],a_{n}]\,.$
@@
When we say that a pro-$p$ group $G$ is finitely generated, or that $X$ is a generating set for $G$, it will always mean generation in the topological sense. We recall that if $w_{0}(G)$ denotes the smallest cardinality of a fundamental system of neighborhoods of $1$ in $G$ – it’s local weight – then $w_{0}(G)=\max\{d(G),|\mathbb{N}|\}$ ([41, Cor. 2.6.3]). In particular, every closed subgroup of a finitely generated pro-$p$ group is at most countably generated.
@@ 2.2. Torsion-free virtually free-by-$\mathbb{Z}_{p}$ groups
There are many abstract torsion-free groups which are virtually free-by-$\mathbb{Z}$ but not free-by-$\mathbb{Z}$ themselves. One explicit family of such groups is given by the presentations $\Gamma_{e}=\langle a,b,x,y\mid[a,b]^{e}=[x,y]^{e}\rangle$ for $e>1$ (see [5, Example 7.2]). More examples arise amongst 3-manifold groups such as fundamental groups of knot complements with Alexander polynomial $1$. One family of such knots is given by the Pretzel links $L(2m+1,2n+1,2q+1)$ satisfying $(m+n+1)(m+p+1)=m(m+1)$, such as L(3,5,7)L(-3,5,7). The condition on the Alexander polynomial implies that $[G,G]$ is perfect, and such groups virtually fibre over $\mathbb{Z}$ by being virtually special as a combined consequence of the works of I. Agol, P. Przytycki and D. Wise([1, 2, 39]).
In the pro-$p$ case we can construct examples of torsion-free pro-$p$ groups which are virtually free-by-$\mathbb{Z}_{p}$ but not free-by-$\mathbb{Z}_{p}$ themselves using the ideas of [5].
@@
We divide the proof into two lemmas. Let $F=F(x_{1},x_{2})$ be another free pro-$p$ group on two generators and consider the surjective homomorphism $\varphi\colon G\to F$ defined by $\varphi(a_{i})=\varphi(b_{i})=x_{i}$ for $i=1,2$.
Lemma 2.2.
@@
Let $V=\Phi(F)$. Then, it is clear that $x_{1},x_{2}]$ is not contained in $[V,V]$ and hence is a non-trivial element of $M=V/[V,V]\simeq\mathbb{Z}_{p}^{n}$. Observe that $x_{1},x_{2}][V,V]$ has finitely many $F$-conjugates which are linearly independent modulo $[V,V]$, and so one can find a homomorphism $\psi_{0}\colon M\to\mathbb{Z}_{p}$ such that $\psi_{0}(x^{-1}[x_{1},x_{2}]x[V,V])=1$ for all $x\in F$. The desired $\psi$ is then the composition of the natural projection $V\to M$ with $\psi_{0}$. ∎
@@
Write $G$ as the proper free pro-$p$ product with amalgamation $G=A\amalg_{C}B$ of the free pro-$p$ groups $A=F(a_{1},a_{2})$, $B=F(b_{1},b_{2})$ and cyclic subgroup $C=\langle[a_{1},a_{2}]^{p}\rangle\simeq\langle[b_{1},b_{2}]^{p}\rangle$ ([41, Exer. 9.2.6a)]). Let $(\mathcal{G},X)$ be the graph of pro-$p$ groups consisting of two vertices $A$ and $B$ and an edge $C$ connecting them such that $\pi_{1}(\mathcal{G},X)\simeq G$, and let $Y$ be the standard $p$-tree associated with $(\mathcal{G},X)$ ([40, Sec. 4]). We recall that:
@@
[…]
@@
Given that $a_{1},a_{2}]=[b_{1},b_{2}]$ is an element of both $U\cap A$ and $U\cap B$ that is not trivial under $\theta$, $V(Z)$ is finite (and in fact consists of two elements).
[…]
@@
and $Z$ is finite. From [42, Thm. 6.6.1], we conclude that the induced map $\pi_{1}(\mathcal{Z},Z)\to K$ is an isomorphism. ∎
Proof of Proposition 2.1.
Again write $G$ as the proper free pro-$p$ product with amalgamation $G=A\amalg_{C}B$ as in Lemma 2.3. Since both $A$ and $B$ are torsion-free, $G$ is also torsion-free by [40, Thm. 4.2(b)]. By [27, Corollary 3.5], $\Gamma$ is residually-$p$, and so it embeds into $G$. Since $a_{1},a_{2}]$ and $b_{1},b_{2}]$ are distinct elements in $\Gamma,\Gamma]\subseteq[G,G]$ with the same $p$-th power, the group $[G,G]$ cannot be free pro-$p$. Hence, $G$ itself cannot be free-by-$\mathbb{Z}_{p}$. However, we’ve shown in Lemma 2.3 that $G$ contains an open subgroup $U$ that is free-by-$\mathbb{Z}_{p}$. ∎